Passengers in PIA flight 123 that is flying west at speed of 900km/h observes F-16 flying for 14August parade rehearsal that passes the PIA plane horizontally 60° southwest direction. Peoples in PIA plane feels that F-16 moving away from them at 70° angle. Determine the true velocity of F-16. 


1
Expert's answer
2020-08-13T12:29:19-0400

As per the given question,

Velocity of the flight 123 "(v)=-900\hat{i}"

Let the velocity of the F-16 is "(v_2)=-v\cos60\hat (i)-v\cos (60)\hat{j}"

observed angle with respect to flight 123"=70^\circ"

Hence, "\tan (70^\circ) =\frac{-v\sin (60^\circ)}{-900-v\cos(60^\circ)}"

"\Rightarrow 2466 -v\cos (60^\circ) =- v\sin 60^\circ"

"\Rightarrow v=\frac{2466}{\cos 60- \sin 60}=6737km/hour"